OpenStudy (anonymous):

What is the sum of the first five terms of a geometric series with a1 = 10 and r = 1/3

OpenStudy (freckles):

Can you write out the first 5 terms?

OpenStudy (freckles):

the first term is already given as 10

OpenStudy (freckles):

what is the second term?

OpenStudy (freckles):

i will help you out these are the first five terms \[a_1,a_1r,a_1r^2,a_1r^3,a_1r^4\]

OpenStudy (freckles):

where a_1=10 and r=1/3

OpenStudy (freckles):

so the sum is \[a_1+a_1r+a_1r^2+a_1r^3+a_1r^4 \\ =a_1(1+r+r^2+r^3+r^4)\]

OpenStudy (freckles):

or you could use the formula for a sum if you so choose

OpenStudy (freckles):

\[a_1 \frac{1-r^5}{1-r}\]

OpenStudy (anonymous):

what about 1/3?

OpenStudy (freckles):

I don't think the sum will be 1/3

OpenStudy (freckles):

replace the a_1's with 10 and replace the r's with 1/3 you can use the formula or just add up those five terms I mentioned

OpenStudy (anonymous):

\[10+10(1/3)+10(1/3)^2+10(1/3)^3+10(1/3)^4+10(1/3)^5\] is this what you mean? #freakles

OpenStudy (freckles):

well that is first 6 terms you are adding

OpenStudy (freckles):

you want to add the first 5 so leave the last one off